OpenStudy (anonymous):

1. Which of the following is the slope between the points (1, 5) and (5, 2)? (4 points) – – 2. Which of the following is the slope between the points (3, –2) and (8, 4)? (4 points) – – 3. Identify the correct x intercept and y intercept of the equation 4x – 3y = 24. (4 points) x intercept = (6, 0); y intercept = (0, –8) x intercept = (–6, 0); y intercept = (0, 8) x intercept = (–8, 0); y intercept = (0, –6) x intercept = (8, 0); y intercept = (0, 6)

OpenStudy (anonymous):

1. The formula for slope is: m = y2 - y1/ x2 - x1 if we plug it in, we get m = 2 - 5/ 5 - 1 which is -3/4 and there's your slope:) 2. Using the formula for slope as I showed you in the first question: m = y2 - y1/ x2 - x1 if we plug it in, we get m= 4-3/8-(-2) which is 1/10 and there's the slope :D 3. For this one, you can either plug them in (since it's multiple choice) or you can solve it as it is. start by solving for y 4x – 3y = 24 subtract 4x from each side to get: -3y = 24 - 4x divide each side by -3 to get: y = -8 + 4/3x OR y = 4/3x - 8 and there is your slope (4/3x), and y-intercept(-8). and now if you look above in the choices they gave you to choose from, find the one with the y-intercept of -8, which is "x intercept = (6, 0); y intercept = (0, –8)" :) hope i helped!

OpenStudy (anonymous):

um. did you get it? (or ...)